Assume that a car weighing 1200kg is traveling east at a velocity of 12.0 m/s. Another car, 40 meters away, weighs 1800 kilograms and travels at a speed of 20.0 meters per second to the east. How do you determine the location of the system's center of mass when it consists of two cars? How do you calculate each car's momentum? What about the overall system's momentum?

1
Expert's answer
2021-12-10T11:30:18-0500

Solution:


(a) "X_C=\frac{m_1x_1+m_2x_2}{m_1+m_2}="


"=\frac{1200\cdot(40+12t)+1800\cdot(20t)}{1200+1800}=16+16.8t" . If "t=0" then "X_C=16\ (m)" .


(b) "p_1=m_1v_1=1200\cdot12=14400\ (kg\cdot m/s)"

"p_2=m_2v_2=1800\cdot20=36000\ (kg\cdot m/s)"


(c) "p=p_1+p_2=14400+36000=50400 \ (kg\cdot m/s)"